# The T2K Collaboration First differential measurement of the single production cross section in neutrino neutral-current scattering K. Abe

### v2

@@
††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain.††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University ”MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
@@ Abstract
Since its first observation in the 1970s, neutrino-induced neutral-current single positive pion production (NC1) has remained an elusive and poorly understood interaction channel. This process is a significant background in neutrino oscillation experiments and studying it further is critical for the physics program of next-generation accelerator-based neutrino oscillation experiments. In this Letter we present the first double-differential cross-section measurement of NC1 interactions using data from the ND280 detector of the T2K experiment collected in -beam mode. We compare the results on a hydrocarbon target to the predictions of several neutrino interaction generators and final-state interaction models. While model predictions agree with the differential results, the data shows a weak preference for a cross-section normalization approximately 30% higher than predicted by most models studied in this Letter.
@@
Figure 1: Expected event rate, in arbitrary units, for CC0 (T2K/Hyper-K) and CC-inclusive (DUNE) muon-neutrino interactions at the Hyper-K (top) and DUNE (bottom) experiments, shown as a function of true neutrino energy. Also overlaid is the expected rate of the NC1 background in each case. The oscillation parameters used are from NuFit5.0 Esteban et al. (2020), and use GENIEv3 Andreopoulos et al. (2010, 2015) to predict the interaction rates using T2K Abe et al. (2013) and DUNE Abi et al. (2020a, b) neutrino fluxes.
@@
Analysis samples. Events with a vertex reconstructed in the fully-active hydrocarbon target FGD1 are classified in four different regions: a signal-enriched sample and three background-enriched samples. The signal definition includes all and interactions without charged leptons in the final state, a single positive pion, any number of protons with momentum below the detection threshold of 200 MeV/c (See details in Ref. Abe et al. (2025)), and no other mesons or other charged particles in the final state. Due to ND280’s acceptance and particle identification capabilities, the measurement is reported in the kinematic Region of Interest (RoI) defined by the conditions 0.5 and GeV/. The neutrino generator NEUT v5.4.0 Hayato (2009a); Hayato and Pickering (2021) is used with T2K’s flux and detector simulation pipeline Abe et al. (2011) to predict the rate of these events. Over 200 signal events are expected overall, with approximately 165 in the signal sample at 30.5% efficiency and 51.4% purity, and an integrated detector uncertainty below 5%. Full selection details and performance metrics are presented in Ref. Abe et al. (2025).
@@
The model independence of the cross-section extraction was validated through extensive tests using simulated datasets with controlled variations. In all cases, small or negligible measurement biases were observed, confirming the robustness of the method.
@@
Interaction Models. We compare the extracted double-differential and integrated cross section to several interaction generators using NUISANCE Stowell et al. (2017): NEUT v5.6.2 Hayato (2002, 2009b); Hayato and Pickering (2021), NuWro v19.02.2 Golan et al. (2012a, b) and GENIEv3 Andreopoulos et al. (2010, 2015) using the AR23_20i configuration. To investigate the role of final-state interaction (FSI), we consider five different FSI models from GENIE Dytman et al. (2021), with all other parameters set to the G18_10_00_000 configuration. These use: the INTRANUKE Dytman (2009) hA single-step empirical model (G18_10a); the INTRANUKE hN full intranuclear cascade (G18_10b); GENIE’s implementation of the INCL cascade Mancusi et al. (2015); Cugnon et al. (2016) (G18_10c); the GEANT4 Bertini cascade Wright and Kelsey (2015) (G18_10d); and an unphysical model in which all final-state interactions have been turned off (G18_10X). AR23_20i is identical to G18_10a for NC pion production, except that the free nucleon pion production tuning from Ref. Tena-Vidal et al. (2021) is applied.
@@
Conclusions. Using the largest data sample of selected NC interactions to date, the double-differential cross section for this channel has been presented for the first time. Overall, the differential cross-section results agree with tested model predictions, though we observe a statistically weak preference for a cross section approximately 30% higher than most model expectations. The observed variations across FSI models exceed those between neutrino interaction generators, emphasizing the need to improve our understanding of FSI.
@@
This work establishes the foundation for more precise measurements of this channel in T2K and Hyper-Kamiokande using the ND280 detector. With the recent upgrade of the ND280 detector Abe et al. (2019) and planned increases in data collection over the coming years, we anticipate continued improvements in understanding this channel, enabling more precise measurements of neutrino oscillation parameters.

### v3

@@
††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University ”MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
@@ A. Bubak
J.A. Caballero
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
F. Cadoux
@@ C. Forza
J. García-Marcos
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ V. Glagolev
R. González Jiménez
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
J. González Rosa
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
@@ A. Holin
M.C.F. Hooft
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ A. Izmaylov
N. Jachowicz
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ A. Konaka
A. Kumar Jha
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ A. Mefodiev
G.D. Megias
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
@@ V.Q. Nguyen
K. Niewczas
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ Abstract
Since its first observation in the 1970s, neutrino-induced neutral-current single positive pion production (NC1) has remained an elusive and poorly understood interaction channel. This process is a significant background in neutrino oscillation experiments and studying it further is critical for the physics program of next-generation accelerator-based neutrino oscillation experiments. In this Letter we present the first double-differential cross-section measurement of NC1 interactions using data from the ND280 detector of the T2K experiment collected in -beam mode. The measured flux-averaged integrated cross-section is . We compare the results on a hydrocarbon target to the predictions of several neutrino interaction generators and final-state interaction models. While model predictions agree with the differential results, the data shows a weak preference for a cross-section normalization approximately 30% higher than predicted by most models studied in this Letter.
@@
Figure 1: Expected event rate, in arbitrary units, for CC0 including and Cherenkov thresholds, (T2K/Hyper-K) and CC-inclusive (DUNE) muon-neutrino interactions at the Hyper-K (top) and DUNE (bottom) experiments, shown as a function of true neutrino energy. Also overlaid is the expected rate of the NC1 background in each case. The oscillation parameters used are from NuFit5.0 Esteban et al. (2020), and use GENIEv3 Andreopoulos et al. (2010, 2015) truth information to predict the interaction rates using T2K Abe et al. (2013) and DUNE Abi et al. (2020a, b) neutrino fluxes.
@@
Analysis samples. Events with a vertex reconstructed in the fully-active hydrocarbon target FGD1 are classified in four different regions: a signal-enriched sample and three background-enriched samples. The signal definition includes all and interactions without charged leptons in the final state, a single positive pion, any number of protons with momentum below the detection threshold of 200 MeV/c (See details in Ref. Abe et al. (2025)), and no other mesons or other charged particles in the final state. Due to ND280’s acceptance and particle identification capabilities, the measurement is reported in the kinematic Region of Interest (RoI) defined by the conditions 0.5 and GeV/. The neutrino generator NEUT v5.4.0 Hayato (2009a); Hayato and Pickering (2021) is used with T2K’s flux and detector simulation pipeline Abe et al. (2011) to predict the rate of these events. Over 200 signal events are expected overall, with approximately 165 in the signal sample at 30.5% efficiency and 51.4% purity, and an integrated detector uncertainty for the predicted event rate below 5%. Three analysis sidebands enriched in background events are used in the analysis. This includes a sample with 81.0% purity in CC events, the most numerous background events in the signal sample corresponding to 28.2% of its selected events. The other sidebands are enriched in events with additional tracks in the neutrino target and events where the main track is identified as a proton instead of a . Full selection details and performance metrics are presented in Ref. Abe et al. (2025).
@@
The model independence of the cross-section extraction was validated through extensive tests using alternative datasets with controlled variations. In all cases, small or negligible measurement biases were observed, confirming the robustness of the method.
@@
Interaction Models. We compare the extracted double-differential and integrated cross section to several interaction generators using NUISANCE Stowell et al. (2017): NEUT v5.6.2 Hayato (2002, 2009b); Hayato and Pickering (2021), NuWro v19.02.2 Golan et al. (2012a, b) and GENIEv3 Andreopoulos et al. (2010, 2015) using the AR23_20i configuration. These represent a range of commonly used models, which differ in a variety of details including hard scattering processes, nuclear modeling and re-interactions of interaction products within the nucleus (Final State Interactions, FSI). Additionally, we investigate the role of FSI using five different GENIE model variations, where the only change is the FSI model used. These represent a broad range of models of varying levels of sophistication ranging from no FSI, and full semi-classical cascade models. They include: the INTRANUKE Dytman (2009) hA single-step empirical model (G18_10a); the INTRANUKE hN full intranuclear cascade (G18_10b); GENIE’s implementation of the INCL cascade Mancusi et al. (2015); Cugnon et al. (2016) (G18_10c); the GEANT4 Bertini cascade Wright and Kelsey (2015) (G18_10d); and an unphysical model in which all final-state interactions have been turned off (G18_10X). AR23_20i uses the hA model and is identical to G18_10a for NC pion production, except that the free nucleon pion production tuning from Ref. Tena-Vidal et al. (2021) is applied.
@@
Conclusions. Using the largest data sample of selected NC interactions to date, the double-differential cross section for this channel has been presented for the first time. Overall, the differential cross-section results agree with tested model predictions, though we observe a statistically weak preference for a cross section approximately 30% higher than most model expectations. The variations observed across FSI models are greater than those between neutrino interaction generators, and could significantly reduce the observed normalization discrepancy. This highlights the critical need to enhance our understanding of final-state interactions. We anticipate that NC predictions will benefit from forthcoming high-statistics neutrino interaction measurements from T2K and other experiments. In particular, future measurements that precisely measure final-state hadrons in interactions both with and without final-state pions and utilizing transverse kinematic imbalance observables will be especially valuable to inform existing nuclear ground state and final-state-interaction models, directly benefiting NC model predictions.
@@
This work establishes the foundation for more precise measurements of this channel in T2K and Hyper-Kamiokande using the ND280 detector. With the recent upgrade of the ND280 detector Abe et al. (2019) and planned increases in data collection over the coming years, we expect continued improvements in the understanding of this channel, enabling more precise measurements of neutrino oscillation parameters. We anticipate that the upgraded ND280 geometry, with an enhanced angular acceptance, should allow to greatly extend the phase space of this measurement in the future to regions where the pion is emitted at large angles with respect to the beam, perhaps even including samples with backwards-going pions. An increase in the signal purity is also expected thanks to the better granularity and timing capabilities of the SuperFGD hydrocarbon target and the addition of Time-of-flight panels in the detector. Unlike the target used in this analysis, SuperFGD is also capable of identifying neutrons, opening new avenues to further understand this channel.
The measurement results have an associated data release that can be found in Ref. T2K Collaboration (2025).